OpenStudy (anonymous):

What is the midpoint of (3, 4) and (–1, –2)?

OpenStudy (anonymous):

Do you know the midpoint formula?

OpenStudy (johnweldon1993):

To find the midpoint...use the fact that \[\frac{ x1 + x2 }{ 2 }\] is your x midpoint and \[\frac{ y1 + y2 }{ 2 }\] is your y midpoint...so (3, 4) and (–1, –2) \[\frac{ 3 - 1 }{ 2 } = x , \frac{ 4 - 2 }{ 2 } = y\]

OpenStudy (anonymous):

(1,1)

OpenStudy (anonymous):

Correct, @FutureMathProfessor ... Tha answer is (1,1)

OpenStudy (anonymous):

Thank you @FutureMathProfessor and @Uchiha_Itachi ;D

OpenStudy (anonymous):

Any time babe :) Medal? :D

OpenStudy (anonymous):

No problem, but any more questions?

OpenStudy (anonymous):

If not, then see you around!